Circular No. 7832

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NON-KREUTZ NEAR-SUN COMET GROUPS It is well-known that some 94 percent of the faint comets observed near

the sun by the SOLWIND, SMM, and SOHO coronagraphs are members of the Kreutz sungrazing group. Although there may be significant differences in the other orbital elements, in several instances pairs of these comets have come to perihelion within only a matter of hours (Sekanina 2002, Ap.J. 566, 577). Less well known is the fact that there are also interrelations among perhaps half of the remaining near-sun SOHO comets. These include the close temporal pairs C/2000 C2–2000 C5, C/2000 C3–2000 C4, and C/2000 Y6–2000 Y7 (IAUC 7364, 7567). The pairs C/1999 J6–1999 U2 (separated by 4.5 months) and C/1997 L2–2001 X8 (separated by 4.5 years) were pointed out on MPEC 2001-C28 by B. G. Marsden, Harvard-Smithsonian Center for Astrophysics, and M. Meyer, Kelkheim, Germany, respectively.

Meyer remarked that his pair is also likely to be associated with C/2001 E1, and more recently he has suggested that C/2001 T1, as well as the C/2000 C2–2000 C5 pair, belong to a six-member group, the ambiguous sense of orbital motion of these poorly observed comets having caused the wrong solutions to be adopted. Marsden adds that the wrong solutions were also probably adopted for the poorly observed C/2000 C3–2000 C4 pair, which with his pair mentioned above would then form a four-member group.

Possible parabolic orbital solutions (equinox 2000.0) for the members of the Meyer and Marsden groups are collected here as follows:

Comet  T (TT) q (AU)  ω  Ω  i  MPC
C/1997 L2 June 10.87 0.0381 57.^{o}30 72.^{o}62  71.^{o}69  35205
C/2000 C2 Feb. 3.86 0.0370 55.44 73.71  71.35
C/2000 C5 Feb. 7.89 0.0358 54.73 65.16  72.22
C/2001 E1 Mar. 15.64 0.0357 58.28 72.24 73.37 44504
C/2001 T1 Oct. 9.17 0.0364 57.41 72.56  72.87
C/2001 X8 Dec. 12.86 0.0371 56.15 74.35  72.28  44505
C/1999 J6 May 11.59 0.0492 22.47 81.69  26.53  39791
C/1999 U2 Oct. 25.23 0.0492 22.22 82.05  27.05  36654
C/2000 C3 Feb. 4.59 0.0487 23.47 81.85  24.97
C/2000 C4 Feb. 5.17 0.0487 23.05 81.95  24.97
2002 February 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green